Can you help me solve this? Tom can mow the lawn in 4 hours. Perry can mow the same lawn in 5 hours. How long would it take both of them, working together, to mow the lawn?
Answer by Boreal(15235)

Tom mows 1/4 of the lawn in 1 hour x/4 of it in x hours
Perry can do 1/5 and x/5
Together, they do the full lawn, 1
so (x/4)+(x/5)=1
multiply by 20 to clear fractions
5x+4x=20
9x=20
x=20/9 hours or 2 2/9 hours
during which, Tom mows 20/36 of the lawn and Perry 20/45
this is 5/9 and 4/9 or 1.